Vegas Meeting notes for November 12, 2013
Present: J. Ford, R. McCullough
, A. Kepley, J. Ray, M. Whitehead, R. Prestage, A. Roshi, D. Werthimer, M. Wagner, G. Jones
- Rationalize build environment for BOFs.
- Glenn has this working. He's building bofs now.
- TCPBorphServer-ARP Spamming.
- Ray finished this last week.
- Fix task pinning to cores.
- Joe says this is ready for testing and deciding which cores are used for what.
- Fixing data drop at the end of scans.
- Done according to Mark. Anish/Amanda expressed concern if there is a 20 second lag at the end of a scan.
- Fix for the switching signal phase problem Joe identified
- Fixed as part of the bof rebuild with the new build environment
- When can we rebuild and engineering-test all of modes 1,2,3 with all fixes in place and the rationalized build environment?
- By the end of the day today both h1k and h16k will be done.
- Status of engineering testing of modes 20-29
- No progress due to the build environment, 1/8 spectrum, and ssg follies.
Any other business?